# Magneto-optical response of Dynes superconductors

### v1

@@ Abstract
We investigate the terahertz conductivity of conventional superconductors in Voigt and Faraday magneto-optical configurations. In the Voigt geometry, an ultrathin superconducting film is fully penetrated by the magnetic field which interacts with the spin, thus modifying the magnitudes of the optical gap and of the density of the condensate. We provide an alternative interpretation of the recent experiments showing the gapless conductivity of a Nb film measured by Lee et al. [1] which describes better their data for magnetic field above 1 T. In the Faraday geometry, we analyze the terahertz conductivity of three NbN films with varying thicknesses using the Maxwell-Garnett model, treating vortices as normal-state inclusions within a superconducting matrix. Moreover, we effectively account for ubiquitous pair-conserving and magnetic-field-dependent pair-breaking disorder scattering processes using the model of Herman and Hlubina [2].
[…]
1 Introduction
@@
Magnetic field has dramatic effects on superconductivity—it induces screening currents and the appearance of a vortex lattice (orbital effect), and it interacts with the electron spin (Zeeman effect). The contributions of these two effects to the optical conductivity depend on the orientation of the applied magnetic field with respect to the sample. In the Faraday geometry, the magnetic field is applied perpendicular to the surface, which results in screening currents suppressing the field inside the superconductors, thus the orbital effect usually dominates. In the Voigt geometry, the magnetic field is directed along the surface. In films with a thickness smaller than the penetration depth, the magnetic field is only weakly suppressed inside the film and the Zeeman effect dominates.
@@
Abrikosov developed a theoretical approach able to account for the pair-breaking effect of paramagnetic impurities [3]. Later, Maki [4] showed that the Abrikosov approach could be used for other pair-breaking agents including external magnetic field. Finally, Skalski, Betdeder-Matibet and Weiss [5] worked out an integral expression for optical conductivity which accounts for pair-breaking effects. Their SBW model was recently used to explain THz spectra of niobium, NbN and Nb0.5Ti0.5N films in the Voigt geometry [1, 6, 7]. Xi et al. [6, 7] observed the suppression of the pair-correlation gap $\Delta$ and of the spectroscopic gap $\Omega_{G}$ in ultrathin films of NbN and Nb0.5Ti0.5N in agreement with the SBW model [5], using the pair-breaking parameter $\alpha$ to account for the effects of the magnetic field. While for NbN, $\alpha$ is linearly proportional to $B$ ($\alpha\sim B$), for Nb0.5Ti0.5N, a square law has been observed ($\alpha\sim B^{2}$).
@@
In similar measurements on an ultrathin niobium film [1], Lee et al. managed to fully suppress the spectroscopic gap at 2.4 T for a temperature of 1.5 K. In the interval 2.4–3.5 T, $\Omega_{G}=0$ but $\Delta$ still persists. For magnetic fields above 3.5 T, the superconductivity is lost. Their THz spectra are well described by the SBW model even in the gapless regime. However, in contrast to the observed data, the SBW model predicts a steep fall in the real part of conductivity near low frequencies for magnetic fields of 1–2 T. Moreover, in both mentioned cases [1, 6], the resulting interpretation of the SBW model is questionable, since it uses orbital effect-originated relation between pair-breaking parameter and the second power of the amplitude of the external magnetic field [8] while in Voigt geometry setup the Zeeman splitting effect usually dominates.
@@
In this paper, we propose a way of interpreting the THz spectra in the magnetic field by replacing the SBW model effectively with the one developed by Herman and Hlubina (HH model) [2] including the heuristic dependence of the pair-breaking parameter on the external magnetic field, in a similar fashion as in Ref. [9]. Notice that the HH model has basically the same number of parameters as the SBW model. This makes them equally computationally challenging and allows for a simple comparison of their results.
@@
In the Faraday geometry, there are two approaches to describing the terahertz conductivity. In the first one, vortex cores are treated as normal-state inclusions inside a superconducting environment and this inhomogeneous system can be described as a homogeneous one with an effective complex conductivity. Maxwell-Garnett theory [10] is the most often used [11, 12] as it correctly captures the topology of isolated vortex cores. In this approach, vortices are static and possible vortex motion and related consequences are thus neglected. In the second approach, the two-fluid model is utilized in various theoretical models [13, 14, 15, 16] to describe the vortex dynamics as the dominant contribution to both the real and the imaginary parts of conductivity. The inherent weakness of the two-fluid model is that it does not account for the presence of the superconducting gap in the spectra, which can be justified for frequencies well below the gap but fails in the vicinity of gap and for even higher frequencies. In this article, we propose possible improvements of Xi’s approach [11], where the complex conductivity of the superconducting matrix is not described by its zero-field value but the HH model is used to account for the presence of the pair-conserving and pair-breaking scattering processes, the latter assumed to increase in the presence of the external magnetic field.
2 Theory
In this section, we shortly introduce the fundamental aspects of the SBW and the HH models.
First, the density of states (DOS) needs to be examined. The BCS theory [17] postulates the oversimplified formula for the DOS
@@ 2 Theory
where $\Re$ stands for the real part of the complex value and the parameter $u=E/\Delta$ equals the energy normalized to the value of the gap where E=0 at Fermi level. For energy within the interval $-\Delta<E<\Delta$ there are no states and the DOS diverges at $E=\pm\Delta$. The Zeeman effect leads to a shift in the DOS for up and down spins, thus their respective DOS are $2\mu_{B}B$ apart, where $B$ is the magnetic field and $\mu_{B}=9.274\times 10^{-24}$ J T-1 denotes the Bohr magneton. However, the spin-orbit scattering leads to mixing the states for spin up and down states. The spin-orbit scattering rate is proportional to the fourth power of the atomic number [18, 19], thus the described shift can be clearly observed only in samples with atoms with a low atomic number such as Al films [20]. Here, we study the response of Nb and NbN films in the magnetic field, thus the DOS for up and down spins are smeared into one broad peak. Further details can be found in the Fulde’s review paper [21].
In the limit of large spin-orbit scattering, the SBW model generalizes the DOS from the BCS theory, Eq. (1), and $u$ is evaluated from the following nonlinear equation:
@@
where $\alpha$ is the Maki’s pair-breaking parameter. Eq. (2) has an analytical solution [22] corresponding to the root of the quartic function. The pair-correlation gap $\Delta$ is both temperature and magnetic field-dependent. In zero magnetic field, its value corresponds to the gap value obtained from the optical measurements $\Omega_{G}$. However, with increasing magnetic field, or more generally with increasing $\alpha$, $\Omega_{G}(B)$ decreases faster than $\Delta(B)$. At a critical value of the magnetic field, the superconductor enters a so-called gapless state where $\Omega_{G}$ becomes zero while $\Delta>0$ thus allowing the persistence of the superconducting state. At even higher values of the magnetic field, $\Delta$ becomes zero and the superconductivity is finally fully suppressed.
@@
On the other hand, the Dynes formula [23] accounts for the broadening of the DOS by phenomenologically introducing the broadening parameter $\Gamma$:
@@
$\mathrm{DOS}=\Re\,\left(\frac{E+i\Gamma}{\sqrt{(E+i\Gamma)^{2}-\Delta^{2}}}\right).$ (3)
@@
The microscopic origin of the Dynes formula was unclear for a long time, but in a recent paper [24], it was shown that the formula given by Eq. (3) is valid in systems with a pair-breaking, magnetic, classical disorder provided that the pair-breaking scattering probability distribution has a Lorentzian distribution with width $\Gamma$ within the Coherent Potential Approximation (CPA). In the presented context, the parameter $\Gamma$ accounts for the influence of the magnetic (spin-flip) scattering. In this way we model the influence of Zeeman splitting and spin-orbit interaction for Voigt and the presence of the screening currents in Faraday geometries in a simple phenomenological way. We will also assume that the effective pair-breaking parameter $\Gamma(B)$ is an increasing function of the external magnetic field, playing thus similar role as $\alpha$ in the SBW model. Moreover, the solution of the CPA equations [24] results in the convenient form of the superconducting gap function $\Delta(E)$ and the wavefunction renormalization $Z(E)$ [2]
@@
where [^1] We choose the square root branch so that the signs of $\mathrm{Re}\{\Omega(E)\}$ and $E$ are the same. $\Omega(E)=\sqrt{\left(E+i\Gamma\right)^{2}-\overline{\Delta}^{2}}$ and $\overline{\Delta}=\overline{\Delta}(T)$ corresponds to the temperature dependent gap [25]. This formulation introduces the pair-breaking $\Gamma$ and the pair-conserving $\Gamma_{s}$ scattering rates. It results in the Dynes tunnelling density of states and fulfils the Anderson theorem [26]. Although the main advantage of the HH model from our perspective is the clear identification of the pair-conserving and pair-breaking scatterings in the superconducting state, it also results in a natural normal state limit $\Delta(E)=0$ and $Z_{N}(E)=(1+i\Gamma_{N}/E)$, where $\Gamma_{N}=\Gamma_{s}+\Gamma$ unifies the effect of the disorder, as expected. In the normal state, the overall scattering rate is related to the scattering time $\tau$ in the Drude model for optical conductivity by $\Gamma_{N}=\hbar/(2\tau)$.
2.1 Optical conductivity utilizing HH model
The optical conductivity of Dynes superconductors is calculated from the integral expression resulting from the Green function approach [2, 24, 27]. Furthermore, its numerical evaluation is as costly as that which makes use of the generalized Mattis-Bardeen formula [28] but, unlike the latter, it allows also for pair-breaking processes.
When discussing the optical response of Dynes superconductor, it is customary to follow Nam [29] and introduce three complex functions: the density of states $n(\omega)$, the density of pairs $p(\omega)$ and a function $\epsilon(\omega)$ with dimension of energy
@@ 2.1 Optical conductivity utilizing HH model
The functions $n(\omega)$ and $p(\omega)$ are obviously linked together, and they satisfy the constraint $n^{2}(\omega)-p^{2}(\omega)=1$. Note that the functions $n_{1}(\omega)$, $p_{2}(\omega)$ and $\epsilon_{2}(\omega)$ are even, whereas the functions $n_{2}(\omega)$, $p_{1}(\omega)$ and $\epsilon_{1}(\omega)$ are odd in $\omega$.
@@
Assuming the local limit of the electromagnetic field response, motivated by the characteristic values of the considered mean free paths $\ell\sim 1\mathrm{nm}$, the resulting optical conductivity can be written as [2, 27]
$\sigma_{s}(\omega)=\frac{iD_{0}}{\omega}\int_{-\infty}^{\infty}d\omega^{\prime}\Big{[}\big{(}f(\omega^{\prime}+\omega)-f(\omega^{\prime})\big{)}H_{1}(\omega^{\prime}+\omega,\omega^{\prime})$
$-f(\omega^{\prime})\big{(}H_{2}(\omega^{\prime}+\omega,\omega^{\prime})+H_{2}^{*}(\omega^{\prime},\omega^{\prime}-\omega)\big{)}\Big{]},$ (8)
where $D_{0}=ne^{2}/m$ is the normal-state Drude weight and we utilized the Fermi-Dirac distribution $f(E)=1/\left(1+e^{E/(k_{B}T)}\right)$. Notice also, that eq. 8 is formulated using units $\hbar=1$. As shown in $T=0$ K limit, the spectroscopically gapless state is present for arbitrary $\Gamma>0$ [30].
3 Results
3.1 Voigt geometry
In the Voigt geometry, the in-plane external magnetic field fully penetrates ultrathin superconducting films allowing us to study the effects of a homogeneous magnetic field on their superconducting properties. Here, we partly revisit the work of Lee et al. [1]. The authors measured the complex conductivity of a $d=58$ nm thick Nb film with a critical temperature of $T_{c}=8$ K, see Fig 1. From the imaginary part of the conductivity, the London penetration depth $\lambda_{L}$ was estimated as 0.1 $\mu$m, thus the film was thin compared to its penetration depth, i.e. $\lambda_{L}\gg d$. The magnetic field averaged over the film thickness $d$ can be approximated as $H_{avg}\approx H_{0}(1-d^{2}/(12\lambda_{L}^{2}))$ which confirms, for the given parameters, a very homogeneous internal magnetic field. Physically, such a Nb film is too thin to develop a sufficiently thick layer of screening currents able to screen the magnetic field.

### v3

# Deeper insight into the terahertz response of conventional superconductors under magnetic field

@@ Abstract
We investigate the terahertz conductivity of conventional superconductors in Voigt and Faraday magneto-optical configurations. First, we review theoretical approaches describing the fundamental processes of suppression of superconductivity in magnetic field and how the in-gap states are filled. In the Voigt geometry, thin superconducting films are fully penetrated by the magnetic field which interacts with the spin, thus modifying the magnitudes of the optical gap and of the density of the condensate. In this configuration, we provide an alternative description of the recent experiments showing the gapless conductivity of a Nb film measured by Lee et al. [Nat. Commun. 14, 2737 (2023)], which better fits their data for magnetic fields above 1 T. In the Faraday geometry, we measured and analyzed the terahertz conductivity of three NbN films with varying thicknesses using the Maxwell-Garnett model, treating vortices as normal-state inclusions within a superconducting matrix. In both geometries, the optical conductivity can be comprehensively described by the model of Herman and Hlubina [Phys. Rev. B 96, 014509 (2017)] involving pair-conserving, and magnetic-field-dependent pair-breaking disorder scattering processes.
pacs:
74.78.Db
I Introduction
@@
Magnetic field has dramatic effects on superconductivity—it induces screening currents and the appearance of a vortex lattice (orbital effect) and it interacts with the electron spin (Zeeman effect). The contributions of these two effects to the terahertz (THz) conductivity depend on the orientation of the applied magnetic field with respect to the sample (see Fig. 1). In the Faraday geometry, the magnetic field is applied perpendicular to the surface, which results in screening currents suppressing the field inside the superconductors, thus the orbital effect usually dominates. In the Voigt geometry, the magnetic field is directed along the surface. In films with a thickness smaller than the penetration depth, the orbital effect becomes weaker with decreasing film thickness. The Zeeman effect thus dominates in ultrathin samples [1].
Figure 1: Magneto-optical geometries. $\vec{B}$: external magnetic field. The THz beam propagates perpendicular to the surface of the sample, with a wave vector $\vec{k}_{\mathrm{THz}}$ and an electric field $\vec{E}_{\mathrm{THz}}$.
@@
Abrikosov developed a theoretical approach able to account for the pair-breaking effect of paramagnetic impurities [2]. Later, Maki [3] showed that the Abrikosov approach could be used for other pair-breaking mechanisms including external magnetic field. Finally, Skalski, Betbeder-Matibet and Weiss (SBW) worked out an integral expression for optical conductivity that accounts for pair-breaking effects [4]. The SBW model was recently used to explain the THz spectra of niobium, NbN and Nb0.5Ti0.5N films in the Voigt geometry [5, 6, 7]. Xi et al. [6, 7] observed the suppression of the pair-correlation gap $\Delta$ and of the spectroscopic gap $\Omega_{G}$ in ultrathin films of NbN and Nb0.5Ti0.5N in agreement with the SBW model [4], using the pair-breaking parameter $\alpha$ to account for the effects of the magnetic field. While for NbN, $\alpha$ is linearly proportional to $B$ ($\alpha\sim B$), for Nb0.5Ti0.5N, a square law has been observed ($\alpha\sim B^{2}$).
@@
In similar measurements on a thin niobium film [5], for a temperature of 1.5 K, Lee et al. observed a complete suppression of the spectroscopic gap at 2.4 T. Between 2.4 and 3.5 T, $\Omega_{G}=0$ while $\Delta$ was still non zero. In magnetic fields above 3.5 T, the superconductivity was lost. These THz spectra were well described by the SBW model even in the gapless regime. However, in contrast to the observed data, the SBW model predicts a steep fall in the real part of conductivity near low frequencies for magnetic fields of 1–2 T.
@@
In this paper, we propose a way of interpreting the THz spectra in magnetic field by effectively replacing the SBW model with the one developed by Herman and Hlubina (HH model) [8] which includes a heuristic dependence of the pair-breaking parameter on the external magnetic field, in a similar fashion as in Ref. [9]. Notice that the HH model has basically the same number of parameters as the SBW model. This makes them equally computationally challenging and allows for a simple comparison of their results. Moreover, the resulting analytical formulation of the HH model using a Green’s function allows for a more straightforward insight into several physical quantities, including the density of states (DOS).
@@
In the Faraday geometry, there are two approaches to describing the THz conductivity. In the first one, vortex cores are treated as normal-state inclusions within a superconducting environment and this inhomogeneous system can be described as a homogeneous one with an effective complex conductivity. Maxwell-Garnett theory [10] is the most commonly used [11, 12] as it correctly captures the topology of isolated vortex cores. In this approach, vortices are static and possible vortex motion and related consequences are thus neglected. In the second approach, the two-fluid model is utilized in various theoretical models [13, 14, 15, 16] to describe the vortex dynamics as the dominant contribution to both the real and the imaginary parts of conductivity. The inherent weakness of the two-fluid model is that it does not account for the presence of the superconducting gap in the spectra, which can be justified for frequencies well below the gap but fails in the vicinity of the gap and for higher frequencies. In this article, we propose improvements of the Xi’s approach [11], where the complex conductivity of the superconducting matrix is not described by its zero-field value but the HH model is used to account for the presence of the pair-conserving and pair-breaking scattering processes, the latter gaining in strength in the presence of external magnetic field.
II Theory
II.1 Fundamental aspects of the SBW and the HH models
First, the DOS needs to be examined. The BCS theory [17] postulates the oversimplified formula
@@ II.1 Fundamental aspects of the SBW and the HH models
where $\Re$ stands for the real part of the complex value and the parameter $u=E/\Delta$ equals the energy normalized to the value of the gap where E=0 at Fermi level. For energy within the interval $-\Delta<E<\Delta$ there are no states and the DOS diverges at $E=\pm\Delta$. The Zeeman effect leads to a shift in the DOS for up and down spins, thus their respective DOS are $2\mu_{B}B$ apart, where $B$ is the magnetic field and $\mu_{B}=9.274\times 10^{-24}$ J T-1 denotes the Bohr magneton. However, spin-orbit scattering leads to mixing the states for spin up and down states. The spin-orbit scattering rate is proportional to the fourth power of the atomic number [18, 19], thus the described shift can be clearly observed only in samples with atoms with a low atomic number such as Al films [1]. Here, we study the response of Nb and NbN films in the magnetic field, thus the DOS for up and down spins are smeared into one broad peak. Further details can be found in the Fulde’s review paper [20].
In the limit of strong spin-orbit scattering, the SBW model generalizes the DOS from the BCS theory, Eq. (1), and $u$ is evaluated from the following nonlinear equation:
@@
where $\alpha$ is the Maki’s pair-breaking parameter. The pair-correlation gap $\Delta$ is both temperature and magnetic field-dependent. In zero magnetic field, its value corresponds to the gap value obtained from the optical measurements $\Omega_{G}$. However, with increasing magnetic field, or more generally with increasing $\alpha$, $\Omega_{G}(B)$ decreases faster than $\Delta(B)$. At a critical value of the magnetic field, the superconductor enters a so-called gapless state where $\Omega_{G}$ becomes zero whereas $\Delta>0$ thus allowing the persistence of the superconducting state. At even higher values of the magnetic field, $\Delta$ becomes zero and the superconductivity is finally fully suppressed.
@@
In contrast, the Dynes formula [21] accounts for the DOS broadening by phenomenologically introducing a broadening parameter $\Gamma$:
@@
$\mathrm{DOS}\equiv N(E)=\Re\,\left(\frac{E+i\Gamma}{\sqrt{(E+i\Gamma)^{2}-\Delta^{2}}}\right).$ (3)
A finite value of $\Gamma$ introduces subgap states at the Fermi level due to pair-breaking effects. The formula has been experimentally verified through tunneling spectroscopy measurements in multiple superconducting systems, notably NbN [22, 12], Al [23], and MoC [24]. Furthermore, this approach has successfully described spectroscopic features observed in photoemission studies of YB6 [25].
@@
The microscopic origin of the Dynes formula was unclear for a long time, but in a recent paper [26], it was shown that the formula given by Eq. (3) is valid in systems with a pair-breaking classical disorder provided that the pair-breaking scattering probability distribution has a Lorentzian distribution with width $\Gamma$ within the Coherent Potential Approximation (CPA).
Figure 2: (a) $N(E)$ from the Dynes model (red curve) as an envelope function of different Yu-Shiba-Rusinov bound states ($\alpha=0.04$). (b): The same Dynes $N(E)$ with important values marked.
In Fig. 2$(a)$ we plot the Dynes DOS together with curves corresponding to a DOS resulting from multiple scattering of the electron on a single magnetic impurity (assuming its low concentration), calculated within the T-matrix approximation [27]. The described approach leads to the Yu-Shiba-Rusinov (YSR) sub-gap states in the superconducting gap of the resulting DOS. The center of the sub-gap states can be tuned by the YSR parameter $\gamma_{Y}$ [27] depending on the coupling constant of the magnetic impurity. For simplicity, the value of the other YSR parameter $\alpha$ which expresses the amount of magnetic impurities, was taken as $\alpha=0.04$ in all cases. Notice that the Dynes DOS can be understood as an envelope function of the YSR impurity bands resulting from the multiple scatterings on magnetic impurities with different coupling strengths.
In Fig. 2$(b)$ we indicate the important parameters of the Dynes DOS. The energy of the Dynes DOS maximum in the coherence peak amounts to
$\tilde{E}_{*}^{2}=1+5\gamma^{2}/3+\sqrt{\left(1+5\gamma^{2}/3\right)^{2}-\left(1+\gamma^{2}\right)^{2}}$
where $\tilde{E}_{*}=E_{*}/\Delta$ and $\gamma=\Gamma/\Delta=N(0)/\sqrt{1-N(0)^{2}}$. Note that if $N(0)$ is determined from tunneling measurements at very low temperatures, one can calculate $\gamma$ and $\tilde{E}_{*}$ which allows for a direct estimate of both the relevant gap $\Delta$ and the pair-breaking parameter $\Gamma$.
Also, another interesting energy scale $\tilde{E}_{1}$, the energy value when $N(\tilde{E}_{1})=1$, can be also analytically expressed in the form
$\tilde{E}_{1}^{2}=1-\frac{\gamma^{2}}{3}+\frac{u}{2^{1/3}3}-\frac{2^{1/3}}{3}\frac{\gamma^{2}\left(15-4\gamma^{2}\right)}{u},$
where
$u=\gamma^{\frac{2}{3}}\Big{[}16\gamma^{4}+153\gamma^{2}-27\\ +3\sqrt{3\left(27+194\gamma^{2}+435\gamma^{4}+288\gamma^{6}\right)}\Big{]}^{1/3}.$
In the presented context, the parameter $\Gamma$ accounts for the influence of the magnetic (spin-flip) scattering induced by the pair-breaking effect of the present magnetic field. In this way, we model the influence of Zeeman splitting, orbital effects, and spin-orbit interactions for the Voigt geometry and the presence of the screening currents in the Faraday geometry in a simple phenomenological approach. We also assume that the effective pair-breaking parameter $\Gamma(B)$ is an increasing function, thus playing a role similar to $\alpha$ in the SBW model. Moreover, the solution of the CPA equations [26] results in convenient forms of the superconducting gap function $\Delta(E)$ and of the wavefunction renormalization $Z(E)$ [8]:
@@
where  [^1] We choose the square root branch so that the signs of $\mathrm{Re}\{\Omega(E)\}$ and $E$ are the same. $\Omega(E)=\sqrt{\left(E+i\Gamma\right)^{2}-\overline{\Delta}^{2}}$ and $\overline{\Delta}=\overline{\Delta}(T)$ corresponds to the temperature-dependent gap [29]. This formulation introduces the pair-breaking $\Gamma$ and the pair-conserving $\Gamma_{s}$ scattering rates. It induces the Dynes tunnelling density of states and fulfills the Anderson theorem [30]. Although the main advantage of the HH model from our perspective is the clear identification of the pair-conserving and pair-breaking scatterings in the superconducting state, it also results in a natural normal-state limit $\Delta(E)=0$ and $Z_{N}(E)=(1+i\Gamma_{N}/E)$, where $\Gamma_{N}=\Gamma_{s}+\Gamma$ expresses the total effect of the disorder, as expected. In the normal state, the overall scattering rate is related to the scattering time $\tau$ in the Drude model for optical conductivity by $\Gamma_{N}=\hbar/(2\tau)$.
II.2 Optical conductivity utilizing HH model
The optical conductivity of conventional superconductors is calculated from the integral expression resulting from the Green function approach [8, 26, 31]. Furthermore, its numerical evaluation is as costly as that which makes use of the generalized Mattis-Bardeen formula [32] but, unlike the latter, it also allows for pair-breaking processes.
When discussing the optical response of conventional superconductors, it is customary to follow Nam [33] and introduce three complex functions: the density of states $n(\omega)$, the density of pairs $p(\omega)$ and a function $\epsilon(\omega)$ with a dimension of energy:
@@ II.2 Optical conductivity utilizing HH model
The functions $n(\omega)$ and $p(\omega)$ are obviously linked together, and they satisfy $n^{2}(\omega)-p^{2}(\omega)=1$. Note that the functions $n_{1}(\omega)$, $p_{2}(\omega)$ and $\epsilon_{2}(\omega)$ are even, whereas $n_{2}(\omega)$, $p_{1}(\omega)$ and $\epsilon_{1}(\omega)$ are odd in $\omega$.
@@
Assuming the local limit of the electromagnetic field response, given the characteristic values of the considered mean free paths $\ell\sim 1\,\mathrm{nm}$, the resulting optical conductivity can be written as [8, 31]
[…]
$-f(x)\big{(}H_{2}(x+\omega,x)+H_{2}^{*}(x,x-\omega)\big{)},$
where $D_{0}=ne^{2}/m$ is the normal-state Drude weight and we used the Fermi-Dirac distribution $f(E)=1/\left(1+e^{E/(k_{B}T)}\right)$. Notice also that Eq. (8) assumes $\hbar=1$. As shown in the limit $T=0$  K, the spectroscopically gapless state is present for arbitrary $\Gamma>0$ [34].
III Results
III.1 Voigt geometry
In the Voigt geometry, the in-plane external magnetic field fully penetrates thin superconducting films, allowing us to study the effects of a homogeneous magnetic field on their superconducting properties. Here, we partly revise the work of Lee et al. [5]. The authors measured the complex conductivity of a $d=58$ nm thick Nb film with a critical temperature of $T_{c}=8$ K, see Fig 3. From the imaginary part, the London penetration depth $\lambda_{L}$ was estimated as 0.1 $\mu$m, thus the film was thin compared with its penetration depth, i.e. $\lambda_{L}\gg d$. The magnetic field averaged over the film thickness $d$ can be approximated by $H_{avg}\approx H_{0}(1-d^{2}/(12\lambda_{L}^{2}))$, which confirms, for the given parameters, a very homogeneous internal magnetic field. In fact, such a Nb film is too thin to develop a sufficiently thick layer of screening currents able to screen the magnetic field.